OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12. A) y = -12 - 4/3 x B) y = -4 - 4/3 x C) y = -12 + 4x D) y = -4 + 4x

OpenStudy (loinducoeur):

3y= -12 - 4x y= -4 - 4/3x So, the answer is A :)